# Improved Performances and Motivation in Intelligent Tutoring Systems: Combining Machine Learning and Learner Choice

### v1

@@ 5 Materials and Methods
Several systematic reviews or studies on ITS[48, 49, 45, 1, 3] efficacy pinpointed methodological limitations of this new empirical field (no control group, no initial group equivalence, no pre- and post-intervention measurements, etc[47]) and the great variability of the ITS designs or of their use making it difficult to identify which of the ITS features and/or which of the conditions of learning context of their use 111for instance, ITS can be used alone or mixed with a specific teacher-based instructional setting are critical for successful personalized learning. So in this section, we describe both the ITS features as well as the experimental protocol. As the ITS system presented here was conceived in the context of a project called ’KidLearn’, we also refer to this ITS as the KidLearn system.
@@ 5.1 Activity Space definition
An activity/exercise $e$ is characterized as a particular combination of parameter values inside an activity space where values were selected for each hierarchical group of parameters involved. All the parameters needed to define an activity are instantiated to produce a unique combination of parameter values. The index $u_{i}$ corresponds a selected value $v_{u_{i}}$, for a parameter $a_{i}$, used to generate an activity. To simplify the notation, $u_{i}$ is noted as a given parameter selected value to generate an exercise to differentiate it with $v_{j}$ which defines any values of a parameter. For a group $H_{x}$ with $m$ parameters, the selection of each parameter value produce a combination leading to a singular instantiation of this group $h_{x}={u_{1},\dots,u_{m}}$. After the selection process, a certain number of groups was instantiated, each producing an activity $e={h_{1},\dots,h_{n_{e}}}$, which groups all parameter values that were selected to produce a unique combination. An activity space groups all possible distinct combinations of parameter values that can define an activity in this space. An illustration of a simple example of an Activity Space with the instantiation of an exercise is shown in figure 7.
@@
To compute an exercise, the values are first selected in the primary group $H_{1}$. Here 2 values are possible for 1 parameter. $v_{1,1,2}$ is selected thus $u_{1,1}=2$ and the first group instantiation gives $h_{1}=v_{1,1,2}$, meaning the exercise type is 2. This value is hierarchically linked to the group $H_{3}$. ($H_{2}$ is not used in this case). Then, values in the group $H_{3}$ are selected. Here, $v_{3,1,3}$ is selected so $h_{3}=v_{3,1,3}$, i.e. a difficulty level of 3. All levels are linked to the group of modalities $H_{4}$. Next, values are selected in the group $H_{4}$. For example $v_{4,1,1}$ and $v_{4,2,2}$ leading to $h_{4}=v_{4,1,1},v_{4,2,2}$ meaning a square shape and a blue colour. So the final exercise resulting from the values selection is: $e=h_{1},h_{3},h_{4}\leftrightarrow e=(v_{1,1,2}),(v_{3,1,3}),(v_{4,1,1},v_{4,2,2})$ This means the exercise is of Type 2, level 3 with square shape and blue modalities. The Activity Space groups all the activities that can be generated in this way.
5.2 ZPDES : a combinaison of Multi-Armed Bandit and Intrinsic Motivation theories to manage Teaching Sequences
To address the challenge of managing activities in an Intelligent Tutoring System, the ZPDES (Zone of Proximal Development and Empirical Success) has been proposed as an implementation of the HMABITS architecture[69]. It relies on state-of-the-art Multi-Armed Bandit techniques (MAB) [32, 33] and exploit tthe empirical estimation of learning progress[70] to manage the Activity Space.
To use a casino analogy, multi-armed bandits describe the problem of finding the slot machine that provides the maximum reward, initially unknown, in a set of many different machines. To find the best machine it is needed to spend money exploring each one before being able to always bet on the best one. This boils down to what is called the “exploration/exploitation” trade-off in machine learning and learning processes generally. Here, these approaches are adapted to ITS where the gambler is replaced by the activity manager, the choice of machine is replaced by a choice of activity parameter values, and the reward is replaced by the student learning progress. It is assumed that activities which are currently estimated to provide a good learning progress must be selected more often as described in section.
@@
Due to the combinatorial explosion of parameter values, only one MAB is not used for each possible combination of parameters values in the activity space but a set of simultaneous MAB is used for each group of parameters. The first alternative of considering a given arm for each activity would increase the number of arms. That would increase the number of parameters and the number of trials required to estimate learning progress and thus the learning time. Also, the approach presented here allows the algorithm to identify which features benefit some students more than others.
@@
For each parameter $a_{i}$ inside a group, the quality of its values is evaluated by a bandit expert $w_{i}$. An expert track the reward provided by each value $v_{j}$ on the last several sampling to compute its quality noted $w_{i}(v_{j})$. At any given time, the value to use for each parameter is sampled according to the probabilities given by:
@@
4:     Instantiate primary group $h_{1}\leftarrow sampleValues(H_{1},W_{1})$
5:     $i\leftarrow 1$
6:     {Recursive sample}
7:     while $h_{i}$ require to instantiate a group $H_{x}$ do
8:         $i\leftarrow x$
9:         $h_{i}\leftarrow sampleValues(H_{x},W_{x})$
10:     end while
11:end procedure
12:return $e=h_{1},...,h_{i}$
[h]
The primary group $H_{1}$ has one parameter $a_{1,1}$, for this parameter, the first values is evaluated to have a higher quality than the second one: $w_{1,1,1}\geq w_{1,1,2}$. There are then more chances for the first value to be sampled. Here, the result of the stochastic sample is that $v_{1,1,1}$ is drawn. $v_{1,1,1}$ is linked hierarchically to $H_{3}$, which is the next to be instantiated. The two first values have the same medium quality (same chance to be drawn), and the last one has a low quality (less chance of being drawn). $v_{3,1,2}$ is drawn, which has a dependency with the group H4. H4 is then instantiated and has three parameters. The first parameter $a_{4,1}$ has its first value evaluated to be more interesting than its second one. The quality for the second parameter is ordered as $w_{4,2,2}\geq w_{4,2,3}\geq w_{4,2,1}$ and for the third parameter $w_{4,3,2}\geq w_{4,3,1}$. The three parameters are sampled simultaneously, and $v_{4,1,1}$, $v_{4,2,3}$ and $v_{4,3,2}$ are drawn. Even though $w_{4,2,2}\geq w_{4,2,3}$, $v_{4,2,3}$ had a chance to be drawn, following the process of exploration. The result of the activity generation is: $e=(v_{1,1,1}),(v_{3,1,2}),(v_{4,1,1},v_{4,2,3},v_{4,3,2})$.
@@
A pure selection, based solely on the previous considerations, would explore all possible activities that could be generated in the activity space from the start of the work process. This would have two drawbacks. First, the type and difficulty of the exercises proposed could change too often and reduce the learners’ motivation and engagement. Second, it might not be possible to explore all activity parameters to estimate the learning progress they are providing. To ensure that learners remain in challenging but possible to achieve areas and to be able to assess the quality of each parameter, a mechanism to limit exploration is introduced. Inspired by the Zone of Proximal Development theory [15] and the concept of Flow [16], a pedagogical expert has the possibility to specify rules that define an evolving set of possible/activated activities, judged relevant for the student. These activities keep the student in the zone of Flow or in the Zone of Proximal Develpment (ZPD) based on his successive results (see Fig. 9).
@@
Following this principle, ordered relations between activity parameter values can be defined, leading to a “graph” governing the activity space which are combined with the set of rules that define and manage the ZPD using activation/deactivation mechanisms. Rules and ordered relations are not always defined for each parameter: there is a distinction between subsets of activity parameters that have a clear difficulty progression, and subsets that don’t. For the example used from Sec 5.1, the difficulty levels have a clear ordering while the modalities don’t. In practice, the management of the ZPD proceeds as follows. For activity parameters with no difficulty level relations between their values, a free exploration is allowed and so all of their values are always active. While for parameters that have a clear progression in difficulty, the values will be activated and deactivated depending on the success rate over all active values.
The following mechanism is proposed to generally manage the ZPD. When the recent learner success rate over all active parameter values $\delta_{i,ZPD}$ reaches a value $\lambda_{ZPD}$, the ZPD is expanded to explore another parameter value $v_{i,j}$ by initializing its expert as : $w_{i}(v_{i,j})=\min w_{i}(v^{ZPD})$. When the recent success rate for a particular value $\delta_{v_{i,j}}$ is higher than a threshold $\lambda_{d}$, this activity can be deactivated and removed from the active list of values. These two threshold allow to configure the general exploration behaviour of the algorithm inside the activity space and is illustrated in figure 10.
Figure 10: ZPDES exploration of an activity space, with $\delta_{ZDP}$ the success rate over all active activities, $\lambda_{ZPD}$ the threshold to expand the ZPD, $\delta_{Ax}$ the success rate for the activity $Ax$, and $\lambda_{a}$ the threshold to reach to deactivate an activity.
The main intuition of this process is that when there are some activities whose difficulty grows, the ZPD will have to grow at the same rate. When activities do not have a clear order of difficulty, or when the order might change from person to person, then it is necessary to allow wider exploration of the activities to accommodate individual differences.
Another kind of mechanism is added to allow a more precise and specific parametrization of the ZPD. Indeed, the algorithm needs to be able to activate and deactivate values when the conditions of exploration for an activity parameter depends on another set of parameters. If the value $v_{g,i,j}$ of parameter $a_{i}$ of group $H_{g}$ requires a certain mastery level of value $v_{x,y,z}$, a threshold $\lambda_{v_{x,y,z}}$ is defined corresponding to the success rate a learner must reach with activities using $v_{x,y,z}$ to activate $v_{g,i,j}$. The requirements can be multiple, meaning a values activation can depend on multiple other values, parameters or group to be activated.
@@
13:         Update ZPD: activate/deactivate $w_{i}$ based on $R^{ZPD}$
14:     end for
15:end while
@@
The activities are structured as a graph into an activity space based on expert knowledge (i.e, the "domain model"). The learning paths are then personalized in two ways; first, by exploring and testing continuously various activities inside the activity space in order to assess their didactic potential for the learner’s progress in real time; second, by exploiting and mainly proposing the activities identified as being the most effective for him/her based on the previous assessment.
The simplicity of the activity space on which the approach is based is a first asset. Indeed, it does not rely on any multidimensional student or domain models, and then only the learner’s information about the estimated learning progression for each activity is required. A second asset of LP approach is to leverage an efficient and simple optimization method consisting of prioritize activities’ parameters identified as yielding significant learning outcomes, i.e. Zone of Promixal Development (ZPD). For that, thanks to our multi-armed bandit algorithm, at each step of the optimization process, one arm is chosen and the resulting payoff is estimated, the objective being to discover dynamically the best arm for each student.
@@
So, various and heterogeneous paths are possible across students as expected for taking into account specific learner’s need at a specific point in time. So, the LP approach empowers the ITS adaptivity for diagnosing learner’s ZPD and for making appropriate adjustments to the specific learners’ needs. Taken together, these two main assets enable to quickly design successful ITS (albeit substantial efforts must be done to parameterize the activity graph on which ZPDES runs) and are supportive to hybrid system combining machine learning and rule-based approaches [75].
5.3 Predefined Sequence
@@
To be able to evaluate ZPDES, a baseline has been built as the form of a Predefined Sequence (Predef). This Predefined Sequence is a simple algorithm inspired by mastery learning strategy [76] and instructional design whose reliability has been validated through several user studies [77] and it does not use any machine learning technology. It consists of a sequences of predefined activities organised by difficulty. When a student is working an activity, he needs to have 3 success out of 4 exercises to pass to the next activity, or else he stays on the same activity. The sequence of activities used in this experiment as been designed by an expert in teaching of mathematics for primary school. The sequence is composed of 27 activities we can divide into 8 groups. Each group corresponds to a type with or without decimal. (M, R, MM and RM with integers only, then M, R, MM and RM with decimals).
5.4 Choice
@@
As explained before, studying the impact of choice as a motivational and learning tool, and decoupling it with the actual pedagogical content (e.g. difficulty of exercises) are interesting. ZCO is an experimental condition where the student has choice over a contextual parameter: the objects presented on the screen. In the money game scenario (see below Sec. 5.5), students compose sums corresponding to the item prices or the change to be given when a customer purchases an item. The choice given to the student is between two different objects, but the activity parameterization is the same. The activity is still selected by a ZPDES algorithm and the choice has no impact on the ZPDES operation. The interface used to implement this experimental condition is shown in figure 11. Only the type of exercise and the objects are presented to simplify the interface and reduce the perturbation of the student to a minimum. The position of the choice icon on the screen is determined randomly to avoid presentation bias.

### v2

@@ 5 Materials and Methods
5.1 Algorithmic definitions
Several systematic reviews or studies on ITS[64, 65, 66, 67, 68] efficacy pinpointed methodological limitations of this new empirical field (no control group, no initial group equivalence, no pre- and post-intervention measurements, etc[69]) and the great variability of the ITS designs or of their use making it difficult to identify which of the ITS features and/or which of the conditions of learning context of their use 222for instance, ITS can be used alone or mixed with a specific teacher-based instructional setting are critical for successful personalized learning. So in this section, we describe both the ITS features as well as the experimental protocol. As the ITS system presented here was conceived in the context of a project called ’KidLearn’, we also refer to this ITS as the KidLearn system.
@@ 5.1.1 Activity Space
An activity/exercise $e$ is characterized as a particular combination of parameter values inside an activity space where values were selected for each hierarchical group of parameters involved. All the parameters needed to define an activity are instantiated to produce a unique combination of parameter values. The index $u_{i}$ corresponds a selected value $v_{u_{i}}$, for a parameter $a_{i}$, used to generate an activity. To simplify the notation, $u_{i}$ is noted as a given parameter selected value to generate an exercise to differentiate it with $v_{j}$ which defines any values of a parameter. For a group $H_{x}$ with $m$ parameters, the selection of each parameter value produce a combination leading to a singular instantiation of this group $h_{x}={u_{1},\dots,u_{m}}$.
After the selection process, a certain number of groups was instantiated, each producing an activity $e={h_{1},\dots,h_{n_{e}}}$, which groups all parameter values that were selected to produce a unique combination. An activity space groups all possible distinct combinations of parameter values that can define an activity in this space. An illustration of a simple example of an Activity Space with the instantiation of an exercise is shown in figure 7.
To compute an exercise, the values are first selected in the primary group $H_{1}$. Here 2 values are possible for 1 parameter. $v_{1,1,2}$ is selected thus $u_{1,1}=2$ and the first group instantiation gives $h_{1}=v_{1,1,2}$, meaning the exercise type is 2. This value is hierarchically linked to the group $H_{3}$. ($H_{2}$ is not used in this case).
Then, values in the group $H_{3}$ are selected. Here, $v_{3,1,3}$ is selected so $h_{3}=v_{3,1,3}$, i.e. a difficulty level of 3. All levels are linked to the group of modalities $H_{4}$. Next, values are selected in the group $H_{4}$. For example $v_{4,1,1}$ and $v_{4,2,2}$ leading to $h_{4}=v_{4,1,1},v_{4,2,2}$ meaning a square shape and a blue colour.
So the final exercise resulting from the values selection is:
$e=h_{1},h_{3},h_{4}\leftrightarrow e=(v_{1,1,2}),(v_{3,1,3}),(v_{4,1,1},v_{4,2,2})$
This means the exercise is of Type 2, level 3 with square shape and blue modalities. The Activity Space groups all the activities that can be generated in this way.
@@
5.1.2 ZPDES : a combinaison of Multi-Armed Bandit and Intrinsic Motivation theories to manage Teaching Sequences
To address the challenge of managing activities in an Intelligent Tutoring System, the ZPDES (Zone of Proximal Development and Empirical Success) has been proposed [77, 73]. It relies on state-of-the-art Multi-Armed Bandit techniques (MAB) [78, 79] and exploit the empirical estimation of learning progress[80] to manage the Activity Space.
To use a casino analogy, multi-armed bandits describe the problem of finding the slot machine that provides the maximum reward, initially unknown, in a set of many different machines. To find the best machine it is needed to spend money exploring each one before being able to always bet on the best one. This boils down to what is called the “exploration/exploitation” trade-off in machine learning and learning processes generally. Here, these approaches are adapted to ITS where the gambler is replaced by the activity manager, the choice of machine is replaced by a choice of activity parameter values, and the reward is replaced by the student learning progress. It is assumed that activities which are currently estimated to provide a good learning progress must be selected more often. Prior work showed that this assumption holds for many classes of problems [81] and is intrinsically motivating for people [82].
@@
A set of simultaneous MAB is used for each group of parameters in the activity space rather than a unique MAB used for each possible combination due to the combinatorial explosion of parameter values. The first alternative of considering a given arm for each activity would increase the number of arms. That would increase the number of parameters and the number of trials required to estimate learning progress and thus the learning time. Also, the approach presented here allows the algorithm to identify which features benefit some students more than others.
@@
For each parameter $a_{i}$ inside a group, the quality of its values is evaluated by a bandit expert $w_{i}$. An expert track the reward provided by each value $v_{j}$ on the last several sampling to compute its quality noted $w_{i}(v_{j})$.
At any given time, the value to use for each parameter is sampled according to the probabilities given by:
@@
4:     $i\leftarrow 1$
5:     Instantiate primary group $h_{1}$:
6:          $h_{1}\leftarrow sampleValues(H_{1},W_{1})$
7:     {Recursive sample}
8:     while $h_{i}$ require to instantiate a group $H_{x}$ do
9:         $i\leftarrow x$
10:         $h_{i}\leftarrow sampleValues(H_{x},W_{x})$
11:     end while
12:end procedure
13:return $e=h_{1},...,h_{i}$
The primary group $H_{1}$ has one parameter $a_{1,1}$, for this parameter, the first values is evaluated to have a higher quality than the second one: $w_{1,1,1}\geq w_{1,1,2}$. There are then more chances for the first value to be sampled. Here, the result of the stochastic sample is that $v_{1,1,1}$ is drawn.
$v_{1,1,1}$ is linked hierarchically to $H_{3}$, which is the next to be instantiated. The two first values have the same medium quality (same chance to be drawn), and the last one has a low quality (less chance of being drawn). $v_{3,1,2}$ is drawn, which has a dependency with the group H4.
H4 is then instantiated and has three parameters. The first parameter $a_{4,1}$ has its first value evaluated to be more interesting than its second one. The quality for the second parameter is ordered as $w_{4,2,2}\geq w_{4,2,3}\geq w_{4,2,1}$ and for the third parameter $w_{4,3,2}\geq w_{4,3,1}$.
The three parameters are sampled simultaneously, and $v_{4,1,1}$, $v_{4,2,3}$ and $v_{4,3,2}$ are drawn. Even though $w_{4,2,2}\geq w_{4,2,3}$, $v_{4,2,3}$ had a chance to be drawn, following the process of exploration. The result of the activity generation is:
$e=(v_{1,1,1}),(v_{3,1,2}),(v_{4,1,1},v_{4,2,3},v_{4,3,2})$.
@@
A pure selection, based solely on the previous considerations, would explore all possible activities that could be generated in the activity space from the start of the work process. This would have two drawbacks. First, the type and difficulty of the exercises proposed could change too often and reduce the learners’ motivation and engagement. Second, it might not be possible to explore all activity parameters to estimate the learning progress they are providing. To ensure that learners remain in challenging but possible to achieve areas and to be able to assess the quality of each parameter, a mechanism to limit exploration is introduced.
Inspired by the Zone of Proximal Development theory [84] and the concept of Flow [85], a pedagogical expert has the possibility to specify rules that define an evolving set of possible/activated activities, judged relevant for the student. These activities keep the student in the zone of Flow or in the Zone of Proximal Develpment (ZPD) based on his successive results (see Fig. 9).
@@
Following this principle, ordered relations between activity parameter values can be defined, leading to a “graph” governing the activity space which are combined with the set of rules that define and manage the ZPD using activation/deactivation mechanisms. Rules and ordered relations are not always defined for each parameter: there is a distinction between subsets of activity parameters that have a clear difficulty progression, and subsets that don’t. For the example used from 5.1.1, the difficulty levels have a clear ordering while the modalities don’t. In practice, the management of the ZPD proceeds as follows. For activity parameters with no difficulty level relations between their values, a free exploration is allowed and so all of their values are always active. While for parameters that have a clear progression in difficulty, the values will be activated and deactivated depending on the success rate over all active values. The following mechanism is proposed to generally manage the ZPD. When the recent learner success rate over all active parameter values $\delta_{i,ZPD}$ reaches a value $\lambda_{ZPD}$, the ZPD is expanded to explore another parameter value $v_{i,j}$ by initializing its expert as : $w_{i}(v_{i,j})=\min w_{i}(v^{ZPD})$.
When the recent success rate for a particular value $\delta_{v_{i,j}}$ is higher than a threshold $\lambda_{d}$, this activity can be deactivated and removed from the active list of values. These two threshold allow to configure the general exploration behaviour of the algorithm inside the activity space and is illustrated in figure 1.
The main intuition of this process is that when there are some activities whose difficulty grows, the ZPD will have to grow at the same rate. When activities do not have a clear order of difficulty, or when the order might change from person to person, then it is necessary to allow wider exploration of the activities to accommodate individual differences. Another kind of mechanism is added to allow a more precise and specific parametrization of the ZPD. Indeed, the algorithm needs to be able to activate and deactivate values when the conditions of exploration for an activity parameter depends on another set of parameters. If the value $v_{g,i,j}$ of parameter $a_{i}$ of group $H_{g}$ requires a certain mastery level of value $v_{x,y,z}$, a threshold $\lambda_{v_{x,y,z}}$ is defined corresponding to the success rate a learner must reach with activities using $v_{x,y,z}$ to activate $v_{g,i,j}$. The requirements can be multiple, meaning a values activation can depend on multiple other values, parameters or group to be activated.
@@
13:         Update ZPD:
14:              activate/deactivate $w_{i}$ based on $R^{ZPD}$
15:     end for
16:end while
@@
The activities are structured as a graph into an activity space based on expert knowledge (i.e, the "domain model"). The learning paths are then personalized in two ways; first, by exploring and testing continuously various activities inside the activity space in order to assess their didactic potential for the learner’s progress in real time; second, by exploiting and mainly proposing the activities identified as being the most effective for him/her based on the previous assessment. The simplicity of the activity space on which the approach is based is a first asset. Indeed, it does not rely on any multidimensional student or domain models, and then only the learner’s information about the estimated learning progression for each activity is required.
@@
A second asset of LP approach is to leverage an efficient and simple optimization method consisting of prioritize activities’ parameters identified as yielding significant learning outcomes, i.e. Zone of Promixal Development (ZPD). For that, thanks to our multi-armed bandit algorithm, at each step of the optimization process, one arm is chosen and the resulting payoff is estimated, the objective being to discover dynamically the best arm for each student. So, various and heterogeneous paths are possible across students as expected for taking into account specific learner’s need at a specific point in time. So, the LP approach empowers the ITS adaptivity for diagnosing learner’s ZPD and for making appropriate adjustments to the specific learners’ needs. Taken together, these two main assets enable to quickly design successful ITS (albeit substantial efforts must be done to parameterize the activity graph on which ZPDES runs) and are supportive to hybrid system combining machine learning and rule-based approaches [88].
5.1.3 Predefined Sequence
@@
To be able to evaluate ZPDES, a baseline has been built as the form of a Predefined Sequence (Predef). This Predefined Sequence is a simple algorithm inspired by mastery learning strategy [89] and based on instructional design whose reliability has been validated through several user studies [90]. It does not use any machine learning technology. It consists of a sequences of predefined activities organised by difficulty. When a student is working an activity, he needs to have 3 success out of 4 exercises to pass to the next activity, or else he stays on the same activity. The sequence of activities used in this experiment as been designed by an expert in teaching of mathematics for primary school. The sequence is composed of 27 activities we can divide into 8 groups. Each group corresponds to a type with or without decimal. (M, R, MM and RM with integers only, then M, R, MM and RM with decimals). Table 2 shows the $27$ successive activities for the students following the parameters defined in section 5.2.
G1.1 G1.2 G1.3 G2.1 G2.2 G2.3 G2.4 G3.1 G3.2 G4.1 G4.2 G4.3 G4.4
Ex Type M M M MM MM MM MM R R RM RM RM RM
Difficulty 1 2 3 1 1 2 2 1 2 1 1 2 2
Cents Not - - - - - - - - - - - - -
Remainder - - - - - - - - - - Int - Int
G5.1 G5.2 G5.3 G5.4 G6.1 G6.2 G6.3 G6.4
Ex Type M M M M MM MM MM MM
Difficulty 4 5 5 6 3 3 4 4
Cents Not x€x x€x x,x€ x,x€ - - - -
Remainder - - - - - Int - Dec
Money Type Real Real Real Real Real Real Real Token
G7.1 G7.2 G7.3 G8.1 G8.2 G8.3 G8.4
Ex Type R R R RM RM RM RM
Difficulty 3 3 4 3 3 4 4
Cents Not x€x x€x x,x€ - - - -
Remainder Int - Int - - Int Dec
Money Type Real Real Real Real Real Real Token
Table 2: Detailed predefined sequence
5.1.4 Choice
@@
As explained before, studying the impact of choice as a motivational and learning tool, and decoupling it with the actual pedagogical content (e.g. difficulty of exercises) are interesting. ZCO is an experimental condition where the student has choice over a contextual parameter: the objects presented on the screen. In the money game scenario (see below 5.2), students compose sums corresponding to the item prices or the change to be given when a customer purchases an item. The choice given to the student is between two different objects, but the activity parameterization is the same. The activity is still selected by a ZPDES algorithm and the choice has no impact on the ZPDES operation. The interface used to implement this experimental condition is shown in figure 10. Only the type of exercise and the objects are presented to simplify the interface and reduce the perturbation of the student to a minimum. The position of the choice icon on the screen is determined randomly to avoid presentation bias.